# PolyFly: Polytopic Optimal Planning for Collision-Free Cable-Suspended Aerial Payload Transportation

### v1

@@ VI Conclusion
This paper presented PolyFly, a global planner for a single robot aerial transportation system that models the robot and environment using a set of polytopes and enables the generation of collision-free trajectories in cluttered scenarios. Our proposed method yielded consistently faster trajectories than a popular state-of-the-art baseline in eight different maze-like environments. Experimental results also confirm the trajectories’ dynamic feasibility and applicability of the proposed solution in real-world settings. Future works will explore how the global trajectories can be efficiently updated upon receiving local obstacle information through onboard sensors. We would also like to leverage the latest advances in accelerating polytopic collision avoidance checks [23] to reduce computation.

### v2

@@ VI Conclusion
This paper presented PolyFly, a global planner for a single robot aerial transportation system that models the robot and environment using a set of polytopes and enables the generation of collision-free trajectories in cluttered scenarios. Our proposed method yielded consistently faster trajectories than a popular state-of-the-art baseline in ten different maze-like environments. Experimental results confirm the trajectories’ dynamic feasibility and the applicability of the proposed solution in real-world settings. One limitation is that PolyFly does not yet run at real-time frequencies or account for onboard obstacle sensing. Future work will address this by efficiently updating trajectories using onboard local obstacle information and leveraging advances in polytopic collision-avoidance checks [22] to reduce computation.